# Separate to Collaborate: Dual-Stream Diffusion Model for Coordinated Piano Hand Motion Synthesis

### v1

@@ Abstract
Automating the synthesis of coordinated bimanual piano performances poses significant challenges, particularly in capturing the intricate choreography between the hands while preserving their distinct kinematic signatures. In this paper, we propose a dual-stream neural framework designed to generate synchronized hand gestures for piano playing from audio input, addressing the critical challenge of modeling both hand independence and coordination. Our framework introduces two key innovations: (i) a decoupled diffusion-based generation framework that independently models each hand’s motion via dual-noise initialization, sampling distinct latent noise for each while leveraging a shared positional condition, and (ii) a Hand-Coordinated Asymmetric Attention (HCAA) mechanism suppresses symmetric (common-mode) noise to highlight asymmetric hand-specific features, while adaptively enhancing inter-hand coordination during denoising. The system operates hierarchically: it first predicts 3D hand positions from audio features and then generates joint angles through position-aware diffusion models, where parallel denoising streams interact via HCAA. Comprehensive evaluations demonstrate that our framework outperforms existing state-of-the-art methods across multiple metrics.
@@ 1 Introduction
Driven by advances in artificial intelligence, generating natural and precise human motion from multimodal inputs has become a prominent and promising field of research. Among various input modalities, audio-driven motion synthesis [1, 56, 50, 13] has gained increasing attention, particularly in applications related to embodied intelligence and virtual reality. As an emerging but impactful technology, it has been widely adopted in areas such as gaming, animation, and webcast, fueling innovation and improving user experiences across these domains [11, 29, 51, 27, 30, 52]. A rising subfield within this area is piano-hand motion generation [59, 45, 7], which has garnered growing research interest in recent years. This technology holds promise for developing AI-powered piano tutors that can provide learners with valuable guidance on fingering and rhythm techniques. However, piano performance requires highly complex and coordinated hand movements that involve precise alignment with musical timing, intricate finger trajectories, spatial hand positioning, and fine inter-hand coordination. Traditional rule-based methods [21, 39] struggle to capture the natural fluidity of such motions, while existing data-driven approaches [37, 62, 7], though more expressive, often suffer from such problems as temporal discontinuity and lack of fine-grained control. Consequently, it is of both theoretical and practical importance to explore methods that can accurately model the relationship between audio and hand motion and synthesize realistic, coordinated gestures for piano performance.
@@ 2 Related Work
3D human motion synthesis aims to generate natural human pose sequences and has significant potential for practical applications [23, 17]. Recent advancements in the field of human motion generation can be divided into two subcategories: autoregressive-based and diffusion-based methods. Autoregressive-based methods [8, 32, 9, 55, 61, 63, 10, 34] capture the temporal dependencies in action sequences through frame-by-frame prediction, thereby enabling the generating of complex action sequences. For example, MotionGPT [15] treats human motion as a language and generates coherent motion sequences using large-scale pre-trained models. However, this kind of method is susceptible to error accumulation during the generation of long sequences, resulting in certain deficiencies in the overall quality and stability of the generated actions.
To address this problem, diffusion-based model [5, 24, 16, 28, 20, 25, 47, 26] generation method was introduced into human motion generation. The diffusion model [12, 42, 54, 53] generates samples by reversing the noise process of the long Markov chain, which can effectively alleviate the error accumulation problem while maintaining motion coherence. Representative works include the Motion Diffusion Model (MDM) [43], which uses a classifier-free guidance mechanism to achieve high-quality motion generation; And Guided Motion Diffusion (GMD) [18], by combining text prompts, reference trajectories, and key positions, achieves controllability and diversity of the generation process. The diffusion model has shown advantages in generating high-quality and continuous motion sequences, and has become an important research direction in the current field of human motion generation. It should be noted that in the subfield of piano-audio-driven hand motion generation, recent studies such as pianomotion10M [7] have attempted to apply diffusion models to hand motion synthesis in piano playing scenarios. However, existing methods generally model bimanual movements as a joint representation. This kind of strong coupling assumption ignores the functional differentiation and asynchrony of the left and right hands during performance in the real world. To address these key issues, we propose a dual-stream decoupled generation framework, which improves the detailed expression and musical fit of the generated hand motions while maintaining the coordination of both hands’ movements by designing independent conditional diffusion models and feature interactive attention mechanisms.
@@ 2.2 Piano Guided Hand Motion Generation
Figure 2: The overall framework of our proposed method. It consists of three main modules: (1) Position Predictor, which predicts hand positions from piano audio; (2) Hand-Coordinated Asymmetric Attention (HCAA), applied to the intermediate U-Net features after a Multi-Head Self-Attention (MHSA) layer, enabling asymmetric feature interaction between the two hands. (3) Decoupled Gesture Generator, which employs two independent diffusion models to generate motions for the left and right hands separately.
@@ 3.1 Problem Formulation
Given a N frames piano audio piece $A=\{a_{1},...,a_{N}\}$, our goal is to generate the corresponding position sequence $P=\{p_{1},...,p_{N}\}$ and hand gesture sequence $\Theta=\{\theta_{1},...,\theta_{N}\}$, where $a_{i}\in\mathbb{R}^{3\times 2}$ and $\theta_{i}\in\mathbb{R}^{J\times 3\times 2}$, with $J$ denoting the total number of hand joints (48 per hand in our implementation). The hand position sequence P represents the 3D coordinates of both the left and right hands at each frame, while the hand gesture sequence $\Theta$ captures the Euler angles of each joint, providing a detailed representation of the dynamic hand movements.
@@ 3.2 Training Pipeline
Audio Encoding. Firstly, the input piano audio $A$ is processed through a pretrained audio feature extractor $\Phi_{a}$ to obtain a compact representation of the audio features $f_{a}$ This process is defined as $f_{a}=\Phi_{a}(A)$, where $f_{a}\in\mathbb{R}^{N\times C}$ and $C$ denotes the feature dimension of $\Phi_{a}$ This representation encapsulates essential auditory features, providing high-quality input for subsequent motion generation.
@@ 3.3 Decoupled Hand Motion Generation
1:Piano audio segment $A$
@@
4:$\Phi_{a}$: Audio Feature Extractor (e.g., HuBERT)
5:$\Phi_{r}^{L},\Phi_{r}^{R}$: Audio Feature Refiner for each hand
6:$\Phi_{A2P}^{L},\Phi_{A2P}^{R}$: Audio-to-Position Predictor for each hand
7:$\Phi_{A2M}^{L},\Phi_{A2M}^{R}$: Audio-to-Motion Generator for each hand
8:$\text{HCAA}(\cdot,\cdot)$: Feature Interaction Module
@@
11:$\alpha_{t},\beta_{t}$: Diffusion schedule
12:function GenerateMotion($A$)
@@
[…]
@@ 3.4 Hand-Coordinated Asymmetric Attention
While the left and right hands have their own independence, they still need to work closely together to maintain the fluency and coordination of the playing. Therefore, we not only use a decoupling strategy to generate separate motions for each hand but also introduce a feature interaction module. Specifically, we introduce a Hand-Coordinated Asymmetric Attention (HCAA) mechanism to enhance information exchange between the left and right hand gesture generators. During inference, the diffusion model gradually removes noise through a denoising process.
Figure 3: The way to compute the parameter $\lambda$.
However, at any given timestep during training or inference, the feature representations still contain certain noise interference, which causes the model to focus on irrelevant contextual information. Aiming to mitigate the impact of the aforementioned noise, we perform differential computation on the feature representations of the left and right hands in the intermediate layers of the diffusion model asymmetrically:
$\begin{split}\operatorname{Attn}_{A}(Q_{1},K_{1},V_{1},Q_{2},K_{2},V_{2})&=\text{softmax}\left(\frac{Q_{1}K_{1}^{T}}{\sqrt{d_{k}}}\right)V_{1}\\ &-\lambda\ \text{softmax}\left(\frac{Q_{2}K_{2}^{T}}{\sqrt{d_{k}}}\right)V_{2},\end{split}$ (6)
where $Q_{1}$, $K_{1}$, $V_{1}$ and $Q_{2}$, $K_{2}$, $V_{2}$ represent the Query, Key, and Value matrices of the intermediate latent features in the left and right hand diffusion models, respectively. $d_{k}$ denotes the dimensionality of the input features. As shown in Fig. 3, we calculate $\lambda$ as follows:
@@
Table 1: Quantitative evaluation of our proposed decoupled hand motion generation method and existing models on PianoMotion10M dataset. Regarding PianoMotion10M, we present its performance results across different architectures.1
@@
S2C-Diffusion (ours) - 0.293 0.242 0.028 0.198 0.309 0.215 0.035 0.229 3.012
Figure 4: Illustration of the qualitative results. We visualize the generated gestures across frames using different methods, where GT represents Ground Truth and PianoMotion denotes PianoMotion10M. The anomalous generated gestures in the figure are highlighted with red boxes. In contrast, the relatively natural generated gestures are highlighted with green boxes. As shown in the examples, our method stands out due to its greater motion realism.
@@ 4.1 Experimental Settings
  • Frechet Inception Distance (FID): FID measures the similarity between predicted motions and real data in the feature space by computing the Frechet distance between their feature vectors. This metric evaluates the distributional difference between generated results and real data.
  • Frechet Gesture Distance (FGD): Unlike FID, FGD is used to measure the difference between the predicted gestures and ground truth for a single hand. This metric helps evaluate the similarity of single-hand gestures.
@@
Implementation Details. We select HuBERT [14] as the audio feature extractor and utilize multiple transformer layers [44] as the audio refiner. Specifically, the HuBERT feature extractor produces features with a dimensionality of 1024, while the audio refiner adheres to a standard Transformer architecture, with both the encoder and decoder comprising four layers each. Each training sample has a duration of 8 seconds with a frame rate of 30 frames per second, yielding a sequence of 240 frames. The feature dimension for both the left and right hand joints is 48. For the diffusion model, we use time steps of $T=1000$ and a linearly increasing variance schedule of $\beta_{1}=1e-4$ to $\beta_{T}=0.02$. Our denoising model adopts a 4-layer U-Net architecture, with each layer having dimensions of 256, 512, 1024, and 2048, respectively. We use an Adam optimizer with a learning rate of $52.5e-$ 5. All experiments were conducted on a single NVIDIA RTX 4090 GPU, with a training duration of approximately 4 days.
@@ 4.2 Experimental Results
Quantitative Results. Table. 1 presents a quantitative comparison between the baseline method and our proposed approach across multiple evaluation metrics. During the evaluation, we strictly adhered to the same dataset split and hyperparameter settings as PianoMotion10M [7] to ensure a fair comparison. The results for EmoTalk [37] and LivelySpeaker [62] were obtained by Gan et al. [7] through the application of their method to the corresponding tasks. Specifically, EmoTalk [37] generates poses directly from audio, while LivelySpeaker [62] utilizes a multi-layer perceptron (MLP)-based diffusion framework for motion generation. The results indicate that our method produces a more realistic and coherent generation of hand gestures, surpassing the baseline in key aspects. Compared to the baseline model, our method achieves a significant reduction in the two primary metrics: FGD and FID. This indicates a more accurate modeling of the correspondence between audio and hand motions, as well as more realistic and natural gesture generation. Additionally, consistent improvements in other metrics (e.g., WGD, PD, and Smooth) further demonstrate that our approach excels not only in overall motion quality but also in spatial precision and smooth temporal transitions.
@@ 4.3 Ablation Studies
In this section, we evaluate each design in our framework to verify its effectiveness. Table. 2 presents the results of our ablation studies on different proposed components, including the effect of decoupled noise, the feature interaction module, the position sharing strategy, and our proposed Hand-Coordinated Asymmetric Attention.
@@ 5 Limitations and Future Work
First, the current model operates solely based on audio input. However, in real-world piano performances, a performer’s hand gestures are influenced by various factors beyond the auditory signal, such as personal playing style, physical expression habits, and even emotional nuance. To address this limitation, future work could explore incorporating multi-modal inputs, such as visual recordings, identity or style embeddings, and body posture cues. By fusing these additional modalities, the model could produce more personalized, expressive, and context-aware motions.
Second, the availability of large-scale paired piano-audio and 3D hand motion datasets is still limited. The scarcity of diverse training data constrains the model’s ability to generalize to various playing styles or complex pieces. Future research may consider constructing or expanding datasets that include a broader range of musical genres, playing techniques, and individual performers. This would enable the model to learn a more nuanced mapping between audio and gesture dynamics.

### v2

@@ Abstract
Automating the synthesis of coordinated bimanual piano performances poses significant challenges, particularly in capturing the intricate choreography between the hands while preserving their distinct kinematic signatures. In this paper, we propose a dual-stream neural framework designed to generate synchronized hand gestures for piano playing from audio input, addressing the critical challenge of modeling both hand independence and coordination. Our framework introduces two key innovations: (i) a decoupled diffusion-based generation framework that independently models each hand’s motion via dual-noise initialization, sampling distinct latent noise for each while leveraging a shared positional condition, and (ii) a Hand-Coordinated Asymmetric Attention (HCAA) mechanism suppresses symmetric (common-mode) noise to highlight asymmetric hand-specific features, while adaptively enhancing inter-hand coordination during denoising. Comprehensive evaluations demonstrate that our framework outperforms existing state-of-the-art methods across multiple metrics. Our project is available at S2C.
@@ 1 Introduction
Driven by advances in artificial intelligence, generating natural and precise human motion from multimodal inputs has become a prominent and promising field of research. Among various input modalities, audio-driven motion synthesis [1, 63, 57, 14] has gained increasing attention, particularly in applications related to embodied intelligence and virtual reality [25, 18, 38]. As an emerging but impactful technology, it has been widely adopted in areas such as gaming, animation, and webcast, fueling innovation and improving user experiences across these domains [12, 31, 58, 29, 32, 39, 59]. A rising subfield within this area is piano-hand motion generation [67, 52, 8], which has garnered growing research interest in recent years. This technology holds promise for developing AI-powered piano tutors that can provide learners with valuable guidance on fingering and rhythm techniques. However, piano performance requires highly complex and coordinated hand movements that involve precise alignment with musical timing, intricate finger trajectories, spatial hand positioning, and fine inter-hand coordination. Traditional rule-based methods [22, 45] struggle to capture the natural fluidity of such motions, while existing data-driven approaches [43, 71, 8], though more expressive, often suffer from such problems as temporal discontinuity and lack of fine-grained control. Consequently, it is of both theoretical and practical importance to explore methods that can accurately model the relationship between audio and hand motion and synthesize realistic, coordinated gestures for piano performance.
@@ 2 Related Work
3D human motion synthesis aims to generate natural human pose sequences and has significant potential for practical applications. Recent advancements in the field of human motion generation can be divided into two subcategories: autoregressive-based and diffusion-based methods. Autoregressive-based methods [9, 34, 10, 62, 69, 5, 72, 11, 37] capture the temporal dependencies in action sequences through frame-by-frame prediction, thereby enabling the generating of complex action sequences. For example, MotionGPT [16] treats human motion as a language and generates coherent motion sequences using large-scale pre-trained models. However, this kind of method is susceptible to error accumulation during the generation of long sequences, resulting in certain deficiencies in the overall quality and stability of the generated actions.
To address this problem, diffusion-based model [6, 36, 26, 17, 30, 21, 27, 54, 28, 40, 35, 70, 24] generation method was introduced into human motion generation. The diffusion model [13, 49, 61, 60] generates samples by reversing the noise process of the long Markov chain, which can effectively alleviate the error accumulation problem while maintaining motion coherence. Representative works include the Motion Diffusion Model (MDM) [50], which uses a classifier-free guidance mechanism to achieve high-quality motion generation; And Guided Motion Diffusion (GMD) [19], by combining text prompts, reference trajectories, and key positions, achieves controllability and diversity of the generation process. The diffusion model has shown advantages in generating high-quality and continuous motion sequences, and has become an important research direction in the current field of human motion generation. It should be noted that in the subfield of piano-audio-driven hand motion generation, recent studies such as pianomotion10M [8] have attempted to apply diffusion models to hand motion synthesis in piano playing scenarios. However, existing methods generally model bimanual movements as a joint representation. This kind of strong coupling assumption ignores the functional differentiation and asynchrony of the left and right hands during performance in the real world. To address these key issues, we propose a dual-stream decoupled generation framework, which improves the detailed expression and musical fit of the generated hand motions while maintaining the coordination of both hands’ movements by designing independent conditional diffusion models and feature interactive attention mechanisms.
@@ 2.2 Piano Guided Hand Motion Generation
Figure 2: The overall framework of our method. It consists of three main modules: (1) Position Predictor, which predicts hand positions from piano audio; (2) Hand-Coordinated Asymmetric Attention (HCAA), applied to the intermediate U-Net features after a Multi-Head Self-Attention (MHSA) layer, enabling asymmetric feature interaction between the two hands. (3) Decoupled Gesture Generator, which employs two diffusion models to generate motions for the left and right hands.
@@ 3.1 Problem Formulation
Given a N frames piano audio piece $A=\{a_{1},...,a_{N}\}$, our goal is to generate the corresponding position sequence $P=\{p_{1},...,p_{N}\}$ and hand gesture sequence $\Theta=\{\theta_{1},...,\theta_{N}\}$, where $p_{i}\in\mathbb{R}^{3\times 2}$ and $\theta_{i}\in\mathbb{R}^{J\times 3\times 2}$, with $J$ denoting the total number of hand joints (48 per hand in our implementation). The hand position sequence P represents the 3D coordinates of both the left and right hands at each frame, while the hand gesture sequence $\Theta$ captures the Euler angles of each joint, providing a detailed representation of the dynamic hand movements.
@@ 3.2 Training Pipeline
Audio Encoding. Firstly, the input piano audio $A$ is processed through a pretrained audio feature extractor $\Phi_{a}$ to obtain a compact representation of the audio features $f_{a}$ . This process is defined as $f_{a}=\Phi_{a}(A)$, where $f_{a}\in\mathbb{R}^{N\times C}$ and $C$ denotes the feature dimension of $\Phi_{a}$ . This representation encapsulates essential auditory features, providing high-quality input for subsequent motion generation.
@@ 3.3 Decoupled Hand Motion Generation
1:Piano audio segment AA
@@
[…]
@@
[…]
12:function GenerateMotion(AA)
@@
23:   $\text{log\_var}^{L}\leftarrow\log\left(\frac{(1-\bar{\alpha}_{t-1})\beta_{t}}{1-\bar{\alpha}_{t}}\right)$
@@ 3.4 Hand-Coordinated Asymmetric Attention
While the left and right hands have their own independence, they still need to work closely together to maintain the fluency and coordination of the playing. Therefore, we not only use a decoupling strategy to generate separate motions for each hand but also introduce a feature interaction module. Inspired by Ye et al. [65], we introduce a Hand-Coordinated Asymmetric Attention (HCAA) mechanism to enhance information exchange between the left and right hand gesture generators. During inference, the diffusion model gradually removes noise through a denoising process. However, at any given timestep during training or inference, the feature representations still contain certain noise interference, which causes the model to focus on irrelevant contextual information. Aiming to mitigate the impact of the aforementioned noise, we perform differential computation on the feature representations of the left and right hands in the intermediate layers of the diffusion model asymmetrically:
$\operatorname{Attn}_{s}=\text{softmax}\left(\frac{Q_{s}K_{s}^{T}}{\sqrt{d_{k}}}\right)V_{s}\\ -\lambda_{s}\ \text{softmax}\left(\frac{Q_{\bar{s}}K_{\bar{s}}^{T}}{\sqrt{d_{k}}}\right)V_{\bar{s}},$ (6)
where $s\in\{\text{left},\text{right}\}$, and $\bar{s}$ denotes the opposite hand of $s$. Here, $Q_{s}$, $K_{s}$, $V_{s}$ are the Query, Key, and Value matrices from the intermediate features of stream $s$, and $d_{k}$ denotes the dimensionality of the input features. $\lambda_{s}$ is the parameter $\lambda$ for stream $s$, which is a learnable coefficient controlling the degree of cross-hand suppression. As shown in Fig. 3, we calculate $\lambda$ as follows:
@@
Table 1: Quantitative evaluation of our decoupled hand motion generation method and existing models on PianoMotion10M.
@@
S2C-Diffusion (ours) MM 2025 0.293 0.242 0.028 0.198 0.309 0.215 0.035 0.229 3.012
Figure 4: Illustration of the qualitative results. We visualize the generated gestures across frames using different methods, where GT represents Ground Truth and PianoMotion denotes PianoMotion10M. The anomalous generated gestures in the figure are highlighted with red boxes. In contrast, the relatively natural generated gestures are highlighted with green boxes.
@@ 4.1 Experimental Settings
  • Frechet Inception Distance (FID): FID measures the similarity between predicted gestures and real data in the latent space by computing the Frechet distance between their latent feature vectors. This metric evaluates the distributional difference between generated results and real data.
  • Frechet Gesture Distance (FGD): Unlike FID, FGD is used to measure the difference between the predicted gestures and ground truth in the MANO[46] parameter space for a single hand. This metric helps evaluate the similarity of single-hand gestures.
@@
Implementation Details. We select HuBERT [15] as the audio feature extractor and utilize multiple transformer layers [51] as the audio refiner. Specifically, the HuBERT feature extractor produces features with a dimensionality of 1024, while the audio refiner adheres to a standard Transformer architecture, with both the encoder and decoder comprising four layers each. Each training sample has a duration of 8 seconds with a frame rate of 30 frames per second, yielding a sequence of 240 frames. The feature dimension for both the left and right hand joints is 48. For the diffusion model, we use time steps of $T=1000$ and a linearly increasing variance schedule of $\beta_{1}=1e-4$ to $\beta_{T}=0.02$. Our denoising model adopts a 4-layer U-Net architecture, with each layer having dimensions of 256, 512, 1024, and 2048, respectively. We use an Adam optimizer with a learning rate of 2.5e52.5e-5. All experiments were conducted on a single NVIDIA RTX 4090 GPU, with a training duration of approximately 4 days.
@@ 4.2 Experimental Results
Quantitative Results. Tab. 1 presents a quantitative comparison between the baseline method and our proposed approach across multiple evaluation metrics. During the evaluation, we strictly adhered to the same dataset split and hyperparameter settings as PianoMotion10M [8] to ensure a fair comparison. The results for EmoTalk [43] and LivelySpeaker [71] were obtained by Gan et al. [8] through the application of their method to the corresponding tasks. Specifically, EmoTalk [43] generates poses directly from audio, while LivelySpeaker [71] utilizes a multi-layer perceptron (MLP)-based diffusion framework for motion generation. The results indicate that our method produces a more realistic and coherent generation of hand gestures, surpassing the baseline in key aspects. Compared to the baseline model, our method achieves a significant reduction in the two primary metrics: FGD and FID. This indicates a more accurate modeling of the correspondence between audio and hand motions, as well as more realistic and natural gesture generation. Additionally, consistent improvements in other metrics (e.g., WGD, PD, and Smooth) further demonstrate that our approach excels not only in overall motion quality but also in spatial precision and smooth temporal transitions.
@@ 4.3 Ablation Studies
In this section, we evaluate each design in our framework to verify its effectiveness. Tab. LABEL:tab:ablation presents the results of our ablation studies on different proposed components, including the effect of decoupled noise, the feature interaction module, the position sharing strategy, and our proposed Hand-Coordinated Asymmetric Attention.
@@ 5 Limitations and Future Work
First, the current model operates solely based on audio input. However, in real-world piano performances, a performer’s hand gestures are influenced by various factors beyond the auditory signal, such as personal playing style, sizes of hands and physical expression habits. To address this limitation, future work could explore incorporating multi-modal inputs, such as visual recordings, identity or style embeddings, and body posture cues. By fusing these additional modalities, the model could produce more personalized, expressive, and context-aware motions.
Second, the availability of large-scale paired piano-audio and 3D hand motion datasets is still limited. The scarcity of diverse training data constrains the model’s ability to generalize to various playing styles or complex pieces. Moreover, existing datasets often lack detailed physical constraints, such as finger-key contact labels or joint-angle limits, which are essential for ensuring physically plausible and musically accurate motions. Future research may consider constructing or expanding datasets that include a broader range of musical genres, playing techniques, individual performers and explicit physical annotations. This would enable the model to learn a more nuanced mapping between audio and gesture dynamics.
@@ 6 Conclusion
Acknowledgement
This work was supported by the STI 2030-Major Projects under Grant 2021ZD0201404.